# Dynamic Speculative Agent Planning

### v2

@@ 1 Introduction
LLM-based agents have rapidly moved from controlled demonstrations to real-world deployments, supporting complex multi-step tasks in domains such as autonomous software engineering (Wu et al., 2023), open-world tool use (Shen et al., 2024), and everyday personal assistance (Ge et al., 2024). Although LLM-based agents have achieved widespread adoption, practitioners consistently report substantial end-to-end latency (Hua et al., 2024; Zhang et al., 2024b ; Zhang et al., 2023; Saha et al., 2024). This latency not only degrades user experience but also limits the feasibility of deploying LLM-based agents in time-sensitive scenarios such as real-time decision support (Cai et al., 2025) or interactive tutoring (Modran et al., 2024; Piro et al., 2024).
@@ 4 Dynamic Speculative Planning with User Controllability
Dynamic Speculative Planning begins with a simple intuition: once $\mathbf{A}$} has produced a few actions, it is often possible to forecast how many additional actions it will generate correctly before $\mathbf{T}$} must intervene. Rather than selecting $k$ heuristically or relying on an expensive offline model, we learn to predict it online while the system is running. After one speculative execution breaks either due to a detected mismatch or when a previously the step is exhausted, the system logs a training pair (partial trajectory, $k$) for training while at the same time utilize the latest checkpoint of the predictor to predict the next $k$. These training pairs are collected to train a small DistilBERT-based regressor whose training cost can be negligible compared with inference cost of a single LLM call. To avoid the training time impacting the speed of the system, the predictor training is done asynchronously: for every newly collected pairs of datapoints, an independent thread fine-tunes the predictor to convergence and then writes the updated checkpoint to disk. The orchestration layer always utilizes the most recent checkpoint, so learning progresses without ever blocking agent planning.
The online, multi-threaded training regimen offers two key advantages: (1) Real-time adaptation: the predictor is continuously updated with the most recent user interactions, keeping it aligned with the current usage distribution. (2) Prompt cost savings: the system begins reducing expenses as soon as new data arrive, without waiting to accumulate and curate a large offline dataset.
4.1 Speculative-Step Control as Value Predictions
@@ 5 Main Experiment
The speculative step predictor uses DistilBERT for its efficiency and expressive capacity in forecasting the speculation step $k$ from a given state. The model is fine-tuned using expectile regression to learn the bias and adaptively adjust its predictions based on the collected data. The model is optimized using AdamW with a learning rate of $1\times 10^{-5}$, batch size 1616. Each training process iterates over 33 epochs on a randomly sampled replay buffer (size 2,5002,\!500). For multi-step return estimation, we adopt $\lambda$-return blending ($\lambda=0.95$) to balance bias-variance trade-offs during reward propagation.
@@ 1. Time Decrease ($\Delta{\text{T}}$)
where NN is the total number of tasks. Positive values indicate time savings.
@@ 2. Cost and Token Increase
Here, $ap^{\text{SP}}_{i}$ and $tp^{\text{SP}}_{i}$ represent the number of prompt tokens for the approximation and target in the SP strategy at task $i$, and $ap^{\text{seq}}_{i}$ and $tp^{\text{seq}}_{i}$ represent the number of prompt tokens for the approximation and target in the sequential planning at task ii. Similarly, $ag^{\text{SP}}_{i}$, $tg^{\text{SP}}_{i}$, $ag^{\text{seq}}_{i}$ and $tg^{\text{seq}}_{i}$ represent the corresponding number of generation tokens. For cost, $\text{cost}_{\text{ap}}$ and $\text{cost}_{\text{tp}}$ are the costs for the approximation and target API for prompt tokens, while $\text{cost}_{\text{ag}}$ and $\text{cost}_{\text{tg}}$ are the corresponding costs for generation tokens.
@@ 5. Concurrency Dynamics
where NN is the number of tasks. This metric reflects how aggressively a system exploits parallelism and whether it respects concurrency constraints under varying speculative policies.
Average Speculative Step ($\overline{K}$) In addition to peak concurrency ($\overline{MC}$), which captures brief spikes of load, we aim to characterize the average concurrency of the planning process over the planning process, as it correlates with sustained system pressure. However, the instantaneous concurrency varies continuously over time. Precisely computing the time-averaged number of concurrent API calls would require fine-grained logs of every call’s start and end time, which is often impractical. To address this, we use the average speculative step $\overline{K}$ as a proxy for the true average concurrency. Let $M$ be the total number of planning episodes, and $k_{i}$ be the speculative steps used in episode ii. We define
@@ 5.3 Tunable trade-off control
Applying a fixed positive offset to the unbiased predicted value of $k$ offers a computationally efficient and conceptually straightforward approach to balancing acceleration and cost efficiency. In OpenAGI benchmark, DSP(offset=11) achieves a latency reduction comparable to DSP($\tau=0.95$). Similarly, DSP(offset=22) achieved 97% acceleration by the optimal fixed-$k$ setting where $k=6$, while saving up to 28% cost. These results demonstrate that simple post-prediction adjustment yields robust performance improvements across both latency and cost metrics.
@@ 6.2 Predictor Accuracy over Online Training
Emprically, training with fully supervised targets (i.e., $\lambda=1,\gamma=1$) produced overly conservative predictors. As shown in Table 9, while the Dyn. (SFT) model achieves similar or even lower token cost savings as our TD-trained variant (Dyn. ($\tau$=0.5)), it consistently underperforms in terms of latency reduction. In most cases, the performance was even slower than a fixed $k=2$ baseline, which is the minimal non-trivial setting. Additionally, the Dyn. (SFT) model achieves lower accuracy across all settings compared to Dyn. ($\tau$=0.5) The experimental findings reveal the limitations of the supervised learning approach in this context. We attribute this to high variance in rollout-derived $k$ labels and the averaging tendency of regression, which causes SFT models to underestimate $k$ in uncertain regions.

### v3

@@ 1 Introduction
LLM‑based agents have rapidly moved from controlled demonstrations to real‑world deployments, supporting complex multi‑step tasks in domains such as autonomous software engineering (Wu et al., 2023), open‑world tool use (Shen et al., 2024), and everyday personal assistance (Ge et al., 2024). Although LLM‑based agents have achieved widespread adoption, practitioners consistently report substantial end‑to‑end latency (Hua et al., 2024; Zhang et al., 2024b ; Zhang et al., 2023; Saha et al., 2024). This latency not only degrades user experience but also limits the feasibility of deploying LLM‑based agents in time‑sensitive scenarios such as real‑time decision support (Cai et al., 2025) or interactive tutoring (Modran et al., 2024; Piro et al., 2024).
@@ 4 Dynamic Speculative Planning with User Controllability
Dynamic Speculative Planning begins with a simple intuition: once $\mathbf{A}$} has produced a few actions, it is often possible to forecast how many additional actions it will generate correctly before $\mathbf{T}$} must intervene. Rather than selecting $k$ heuristically or relying on an expensive offline model, we learn to predict it online while the system is running. After one speculative execution breaks either due to a detected mismatch or when a previously the step is exhausted, the system logs a training pair (partial trajectory, $k$) for training while at the same time utilize the latest checkpoint of the predictor to predict the next $k$. These training pairs are collected to train a small DistilBERT-based regressor whose training cost can be negligible compared with inference cost of a single LLM call. To avoid the training time impacting the speed of the system, the predictor training is done asynchronously: for every newly collected pairs of datapoints, an independent thread fine‑tunes the predictor to convergence and then writes the updated checkpoint to disk. The orchestration layer always utilizes the most recent checkpoint, so learning progresses without ever blocking agent planning.
The online, multi-threaded training regimen offers two key advantages: (1) Real‑time adaptation: the predictor is continuously updated with the most recent user interactions, keeping it aligned with the current usage distribution. (2) Prompt cost savings: the system begins reducing expenses as soon as new data arrive, without waiting to accumulate and curate a large offline dataset.
4.1 Speculative‑Step Control as Value Predictions
@@ 5 Main Experiment
The speculative step predictor uses DistilBERT for its efficiency and expressive capacity in forecasting the speculation step $k$ from a given state. The model is fine-tuned using expectile regression to learn the bias and adaptively adjust its predictions based on the collected data. The model is optimized using AdamW with a learning rate of $1\times 10^{-5}$, batch size $16$. Each training process iterates over $3$ epochs on a randomly sampled replay buffer (size 2,5002,\!500). For multi-step return estimation, we adopt $\lambda$-return blending ($\lambda=0.95$) to balance bias-variance trade-offs during reward propagation.
@@ 1. Time Decrease ($\Delta{\text{T}}$)
where $N$ is the total number of tasks. Positive values indicate time savings.
@@ 2. Cost and Token Increase
Here, $ap^{\text{SP}}_{i}$ and $tp^{\text{SP}}_{i}$ represent the number of prompt tokens for the approximation and target in the SP strategy at task $i$, and $ap^{\text{seq}}_{i}$ and $tp^{\text{seq}}_{i}$ represent the number of prompt tokens for the approximation and target in the sequential planning at task $i$. Similarly, $ag^{\text{SP}}_{i}$, $tg^{\text{SP}}_{i}$, $ag^{\text{seq}}_{i}$ and $tg^{\text{seq}}_{i}$ represent the corresponding number of generation tokens. For cost, $\text{cost}_{\text{ap}}$ and $\text{cost}_{\text{tp}}$ are the costs for the approximation and target API for prompt tokens, while $\text{cost}_{\text{ag}}$ and $\text{cost}_{\text{tg}}$ are the corresponding costs for generation tokens.
@@ 5. Concurrency Dynamics
where $N$ is the number of tasks. This metric reflects how aggressively a system exploits parallelism and whether it respects concurrency constraints under varying speculative policies.
Average Speculative Step ($\overline{K}$) In addition to peak concurrency ($\overline{MC}$), which captures brief spikes of load, we aim to characterize the average concurrency of the planning process over the planning process, as it correlates with sustained system pressure. However, the instantaneous concurrency varies continuously over time. Precisely computing the time-averaged number of concurrent API calls would require fine-grained logs of every call’s start and end time, which is often impractical. To address this, we use the average speculative step $\overline{K}$ as a proxy for the true average concurrency. Let $M$ be the total number of planning episodes, and $k_{i}$ be the speculative steps used in episode $i$. We define
@@ 5.3 Tunable trade-off control
Applying a fixed positive offset to the unbiased predicted value of $k$ offers a computationally efficient and conceptually straightforward approach to balancing acceleration and cost efficiency. In OpenAGI benchmark, DSP(offset=$1$) achieves a latency reduction comparable to DSP($\tau=0.95$). Similarly, DSP(offset=$2$) achieved 97% acceleration by the optimal fixed-$k$ setting where $k=6$, while saving up to 28% cost. These results demonstrate that simple post-prediction adjustment yields robust performance improvements across both latency and cost metrics.
@@ 6.2 Predictor Accuracy over Online Training
Emprically, training with fully supervised targets (i.e., $\lambda=1,\gamma=1$) produced overly conservative predictors. As shown in Table 9, while the Dyn. (SFT) model achieves similar or even lower token cost savings as our TD-trained variant (Dyn. ($\tau$=0.5)), it consistently underperforms in terms of latency reduction. In most cases, the performance was even slower than a fixed $k=2$ baseline, which is the minimal non‑trivial setting. Additionally, the Dyn. (SFT) model achieves lower accuracy across all settings compared to Dyn. ($\tau$=0.5) The experimental findings reveal the limitations of the supervised learning approach in this context. We attribute this to high variance in rollout-derived $k$ labels and the averaging tendency of regression, which causes SFT models to underestimate $k$ in uncertain regions.